OpenStudy (anonymous):

Multiply. (x + 6)(x − 6)

OpenStudy (anonymous):

It should be x^2 right

OpenStudy (lilg132):

x * x = x^2 x * -6 = -6x 6 * x = 6x 6 * -6 = -36 so this would be x^2 - 6x + 6x - 36 which is now x^2 - 36

OpenStudy (anonymous):

(a+b)(a-b)=a^2-b^2 (x+6)(x-6)=x^2-36

OpenStudy (anonymous):

Okay thank you guys for your help today it was all very helpful.

OpenStudy (lilg132):

your welcome